OpenStudy (anonymous):

Write as a single fraction. 2+ 5c+9d/6c - 3c-3d/4c

OpenStudy (zehanz):

Let's see...maybe we could simplify 5c - 3c and write everything as a fraction:\[2+5c+\frac{ 9d }{ 6c }-3c-\frac{ 3d }{ 4c }=\]\[\frac{ 2+5c-3c }{ 1 }+\frac{ 9d }{ 6c }-\frac{ 3d }{ 4c }=\]\[\frac{ 2+2c }{ 1 }+\frac{ 9d }{ 6c }-\frac{ 3d }{ 4c }\]Now think about what will be the denominator of the single fraction.

OpenStudy (zehanz):

Oh, and why not first replace the \[\frac{ 9d }{ 6c }\]by (divide by 3)\[\frac{ 3d }{ 2c }\]

OpenStudy (zehanz):

Do you understand that your number is\[\frac{ 2+2c }{ 1 }+\frac{ 3d }{ 2c }-\frac{ 3d }{ 4c }\]